# Future Trends in the Design of Memetic Algorithms: the Case of the Linear Ordering Problem

### v1

@@ Abstract
The way heuristic optimizers are designed has evolved over the decades, as computing power has increased. Initially, trajectory metaheuristics used to shape the state of the art in many problems, whereas today, population-based mechanisms tend to be more effective. Such has been the case for the Linear Ordering Problem (LOP), a field in which strategies such as Iterated Local Search and Variable Neighborhood Search led the way during the 1990s, but which have now been surpassed by evolutionary and memetic schemes. This paper focuses on understanding how the design of LOP optimizers will change in the future, as computing power continues to increase, yielding two main contributions. On the one hand, a metaheuristic was designed that is capable of effectively exploiting a large amount of computational resources, specifically, computing power equivalent to what a recent core can output during runs lasting over four months. Our analysis of this aspect relied on parallelization, and allowed us to conclude that as the power of the computational resources increases, it will be necessary to boost the capacities of the intensification methods applied in the memetic algorithms to keep the population from stagnating. And on the other, the best-known results for today’s most challenging set of instances (xLOLIB2) were significantly outperformed. Instances with sizes ranging from 300 to 1000 were analyzed, and new bounds were established that provide a frame of reference for future research.
@@
An effective decision-making process usually involves solving highly complex optimization problems [1]. The most general classification of optimization techniques distinguishes between exact [2] and approximate [3] optimization. Given the complexity of many real-life optimization problems, many of which are NP-hard, exact optimization methods are usually not applicable within reasonable timeframes [4]. Therefore, approximate methods that strike a balance between the quality of the solution achieved and the computational cost have gained considerable popularity [5]. In this area, metaheuristics offer the possibility of increasing the generality of the developments, thus reducing the effort needed to adapt optimizers to new optimization problems.
@@
One highly complex problem that has benefited greatly from the use of metaheuristics is the Linear Ordering Problem (LOP), which is an NP-hard combinatorial optimization problem with a large number of applications. The LOP was introduced in 1936 by Leontief [6] and formalized in 1958 by Chenery and Watanabe [7]. The simplest definition of it is based on the matrix triangulation problem. Given a matrix , the problem is to determine a simultaneous permutation of the rows and columns of such that the sum of the entries above the main diagonal is maximized [8]. As a result, the problem can be formulated as the search for a permutation that maximizes the equation:
@@
Although the LOP originated in the field of economics, it has been applied in numerous areas, such as: determination of ancestry relationships in archaeology [9], straightline crossing minimization [10], and the development of probabilistic decision-making models [11]. Due to its broad applicability, the LOP has received considerable attention from the scientific community in recent decades [12, 13]. Moreover, since many of these applications involve the use of large matrices, highly complex test sets have been defined. In particular, the most challenging validation set today is the LOLIB library [14], the latest version of which includes instances up to size 1000.
@@
Many diverse algorithms have been proposed to deal with the LOP [8]. This is because, in general, the design of metaheuristics is a complex undertaking, since most of them rely on stochastic processes that are difficult to model and analyze theoretically. Because of this, theoretical analyses often resort to simplifications, meaning they have a restricted scope and the design of optimizers cannot be based solely on theoretical advances [15]. This means that when designing new metaheuristics, researchers must draw on the practical experience gained from similar problems, since they do not have a specific guide of steps to follow [16]. In addition, the relative efficiency between metaheuristics depends largely on the computational capabilities available and the complexity of the instance to be solved, which accounts for the need for a wide variety of proposals [17]. All this explains why the leading metaheuristics designed a few decades ago are very different from the more successful metaheuristics designed in more recent times.
The LOP is a fairly clear example of how computational resources and the complexity of instances affect the design of optimizers. Thus, based on the historical review presented in [8], it is clear that effective optimizers have been evolving from ad-hoc heuristic methods, to trajectory metaheuristics, and on to the current population-based metaheuristics [18, 19].
@@
This paper falls within the field of approximate optimization and aims to contribute to the advancement of the design and development of efficient and robust optimizers. It focuses on the case of the LOP and, in particular, seeks to answer the following research questions:
@@
Are the methods that generate the best solutions for the LOP able to do so in reasonable times, for example, under a day in the most challenging instances currently in use?
@@
Will today’s cutting-edge methods in the LOP continue to retain their prominence in the computing systems of the future, that is, after a significant increase in computing capabilities? Or, on the contrary, will they be unable to scale properly and require methodological changes?
@@
Have optimal solutions already been achieved for the current most challenging datasets, or, on the contrary, is there still room for improvement in the design of the solvers for the LOP?
In order to answer these questions at this time, it is necessary to make use of long-term executions and parallelizations to simulate the increased availability of computational resources. With this goal in mind, we conducted a long-term performance analysis of the current leading method for the LOP, the MA-EDM algorithm [19]. This analysis allowed us to identify some shortcomings in this method, and led us to conclude that, as is usually the case, once the computing power is increased, the algorithms have to be redesigned in order to leverage the additional computing power. In light of the weaknesses identified in the MA-EDM, a new optimizer called Memetic Algorithm with Explicit Diversity Management and Enhanced Intensification (MA-EDM) is proposed that applies an Iterated Local Search (ILS) as an improvement mechanism. As demonstrated in the experimental study, in high-performance environments, MA-EDM vastly outperforms the current best strategy. Moreover, by parallelizing MA-EDM, we were able to simulate scenarios with computing capabilities equivalent to those of a state-of-the-art core in executions lasting over 4 months. This experiment revealed that the best solutions found so far exhibit a significant gap with respect to the optimal ones, indicating that this set of instances has not yet been optimally solved, meaning, therefore, that more methodological advances are required. As an additional contribution, it should be noted that the new Best-Known Solutions (BKS) obtained for instances ranging in size from 300 to 1,000 will provide a frame of reference that will allow other researchers to better evaluate future advances.
The rest of the document is organized as follows. Section 2 discusses the design of approximate algorithms, with a special emphasis on Memetic Algorithms and their adaptations to the LOP. Section 3 focuses on describing the use of high-performance computing environments to achieve the objectives of this research. Section 4 justifies and presents a new algorithmic proposal for the LOP called Memetic Algorithm with Explicit Diversity Management and Enhanced Intensification (MA-EDM). The experimental validation is presented in Section 5. Finally, the last section discusses the conclusions of this research, as well as some potential future areas of work.
@@ 2 Future and Limitations of Memetic Algorithms for the LOP
Evolutionary Algorithms (EAs) are a class of metaheuristics that draw their inspiration from biological evolution [20, 21]. Experience has shown that EAs are very effective at solving difficult optimization problems in a wide variety of fields [22]. However, they typically become more effective when combined with trajectory techniques, or some other refinement mechanisms [23]. This combination of EAs with additional schemes is known as Memetic Algorithms (MAs) [24].
@@
Multiple strategies have been designed to integrate EAs with other mechanisms. In first-generation MAs, enhancement mechanisms were used statically to refine the solutions found by genetic operators. Subsequently, a variety of more complex integration strategies were developed, ranging from schemes that consider a set of improvement mechanisms that are applied adaptively, to schemes that rely on the feedback provided by the improvement mechanisms to adapt internal parts of the EA, such as its fitness function or genetic operators. Despite the advances offered by recent MAs [25], their design, use and implementation is very complex due to the number of parameters and components they consider, as well as to the complex interactions that occur between them. In many problems, the first-generation MAs are still very effective [26, 27, 18], so both the simplest and the most complex methodologies continue to be investigated.
Considering the number of problems in which MAs are the leading strategy [21], and that this is the case with the LOP, this paper focuses on these types of strategies. In particular, in the case of the LOP, the MA that has yielded the current BKS is a first-generation algorithm [19], so this is the method that will be used to conduct the proposed analyses.
The rationale behind the development of MAs is that EAs are often appropriate for exploring different regions and identifying promising areas, but their intensification capabilities are more limited. Because of this, the idea was to integrate EAs with optimizers that had good intensification capabilities. In this sense, hill climbing local search is one of the simplest schemes to intensify a region of the search space. Since, for many problems, finding a local optimum is a computationally very expensive process, in such cases it is not feasible to use more complex schemes, which is why local search remains one of the most used methods in MAs; however, advances in hardware have also made it quite common for more complex trajectory metaheuristics, such as Tabu Search (TS) or Variable Neighborhood Search (VNS), to be considered, and even mathematical programming strategies that are usually even more computationally expensive.
In relation to the design of MAs, the problem of how often and how the improvement mechanism is used is a fundamental task that has been addressed in the literature in various ways. In the context of continuous optimization, the frequency and application point of local searches have been analyzed [28]. This analysis was expanded in [29] for combinatorial optimization problems.
Another crucial point is the definition of the neighborhood for the local search. In order to aid in the selection of the neighborhood, some heuristic procedures for analyzing the fitness landscape have been devised [24]. An in-depth study of neighborhood selection has been proposed, as well as a taxonomy that can be used to classify MAs, taking into account not only the way in which local search is applied, but also the interactions that arise between the different components [30]. It is also important to note that in addition to the design issues implicit to MAs, as with any population-based algorithm, there are several additional characteristics that play an important role in the success of the optimization process. In particular, proper diversity management plays a critical role [31, 32]. Diversity is a property of a group of individuals that quantifies how much they differ from one another. Although there are a variety of principles that can be followed when designing EAs, in the authors’ opinion, a population-based algorithm should ideally work in conditions of high diversity during the initial stages of optimization, and then progressively lose diversity in the vicinity of the global attraction basin. This position has been evaluated for a variety of combinatorial optimization problems. By following this design principle, algorithms have been developed that currently provide the BKS for different optimization problems, such as the Job Shop Scheduling Problem [33] and the Frequency Assignment Problem [34]. The ideal behavior discussed above hardly occurs in practice in EAs, since in most cases, the algorithm tends to converge prematurely towards a suboptimal solution, or to stagnate [32].
@@
In the case of MAs, problems with diversity management appear frequently. For example, in the RHGA algorithm [25], one of the most effective strategies for the TSP, great care had to be taken in selecting the frequency of application of the LKH improvement strategy to avoid premature convergence. This type of problem arises especially in the case of high dimensionality problems [35], so as the instances grow, it is quite common to have to change some design decisions. Issues related to instances of high dimensionality are not unique to MAs [36]. This problem is known as the “curse of dimensionality”, a term proposed by Bellman in 1957 [37], and also appears in other types of optimizers and in machine learning. The curse of dimensionality occurs when an increase in search space size causes a significant loss of performance in optimization algorithms. The increase in the dimensionality or size of the instance causes the search space to grow exponentially in most problems, and as a result, the percentage of solutions evaluated decreases. However, this is not its only drawback, since it can, for example, cause premature convergence in subsets of variables [35]. Therefore, increasing the amount of computing power available (even exponentially) may not solve the challenges that arise when dealing with large instances. To address this problem, various techniques, such as reducing dimensionality and coevolution, have been proposed [38, 39]. This paper takes advantage of the demands associated with high dimensionality problems to fully leverage the metaheuristic schemes known to perform well with the LOP. To do this, we will work with instances of up to , that is, with a search space with approximately candidate solutions, and we will give the algorithm what it supposedly needs: more computing power. It is not at all foolhardy to think that in the near future, computing capacity will advance significantly, which will make more computing power available to algorithms. However, what we do not know is whether today’s methods will be able to continue to perform well when these resources are made available to them. The experiments conducted provide an answer to this question.
@@ 3 Using High Performance Computing Environments
High performance computing (HPC) laboratories consist of a set of multi-core machines that can communicate over high-bandwidth low-latency networks. These laboratories are typically used to reduce the time required to solve highly complex problems. Since the programming and bottlenecks of parallel algorithms differ from those of sequential algorithms, it is quite common for parallel algorithms to be developed that are significantly different from the direct parallelizations of the best sequential algorithms. Thus, in the case of Memetic Algorithms and population-based optimizers in general, various techniques have emerged to take advantage of this type of resource [40], such as master-worker parallelizations, island models, coevolution and cellular models. Among them, the only strategy that maintains the panmictic population-based model of the sequential algorithm is the master-worker model, while the others include changes in the evolutionary process that facilitate the parallelization and treatment of possible bottlenecks.
@@
The objective of using HPC laboratories in this research was not to take advantage of parallel computing resources to obtain high-quality results quickly, but to understand how the effectiveness of one of the leading algorithms in the area of LOP is affected by the increase in computing capabilities, as well as to propose modifications that allow the increase in computing power to be better utilized. Thus, the purpose of this work is not to innovate in parallelization models or find the best parallel models for the LOP, but to generate new knowledge in relation to the optimization of the LOP so as to understand what type of research might be most useful to continue advancing in this field.
As mentioned, dimensionality is one of the characteristics that correlates highly with the difficulty of solving a problem. The case of the LOP is no exception, and over the years, new sets of instances have had to be created that pose a challenge to the best existing algorithms [41]. Although various characteristics have been identified in the instances and the search landscape that affect the difficulty of the solution, the size of the matrix is the most important and easily modifiable [42]. The eXtended LOLIB library (xLOLIB) [42] was proposed for the purpose of creating instances that are complex and exhibit characteristics similar to the matrices that arise in practical problems. Over the years, it has been updated with new instances, and the xLOLIB2 subgroup currently contains instances ranging from 300 to 1000. In recent years, the Best-Known Solutions (BKS) for the largest instances have continued to be updated [43, 19], which is an indication that there is still room for improvement for LOP optimizers when dealing with instances of these sizes. Specifically, MA-EDM [19] is the algorithm that has generated all the BKS for this subgroup of instances, so it is the algorithm that will be used in the analyses.
Considering that one of the purposes of this research is to understand how the MA-EDM and its extensions will behave when using higher performance processors, models should be used that keep the evolutionary process intact. Because of this, the synchronous master-worker model was selected. This scheme is standard in the parallelization of population-based algorithms, and can be programmed using what is commonly known as an embarrassingly parallel model, since the only extension needed is the redistribution of a set of independent tasks among a set of processes that are exclusively dedicated to executing those tasks. In order to facilitate the reproducibility of this research, this section details how to perform the parallelization.
@@
The MA-EDM sequential algorithm is a standard first-generation MA. Its general behavior is described in Algorithm 1. A detailed description of each of its components is available at [19]. Each iteration operates with a population of individuals and consists of a binary tournament-based parent selection phase, a child-reproduction phase that considers the Cycle Crossover (CX), an improvement phase that applies a local search, and a replacement phase that uses the Best-Non Penalized strategy (BNP). Of all the above steps, the local search is the most computationally demanding process, since the neighborhood is considered based on shifts [44], so for instances of size , the step of ensuring that a local optimum has been found using incremental evaluation has a computational cost of — and this is only one part of the computation that is required in the local search — while the rest of the components depend linearly or sublinearly on . Therefore, with large instances like the ones considered in this research, it is natural for the improvement process to be the task that is distributed in the parallelization. Finally, and as mentioned, this distribution is done synchronously so as not to change the evolutionary process of the sequential scheme; that is, once a local search distribution point (or another improvement procedure) is reached, the master process halts until all the local searches are complete.
@@
The flowchart shown in Figure 1 illustrates the operation of the parallel version of MA-EDM, an algorithm that is called Parallel MA-EDM (PMA-EDM). The diagram shows an example with a population of size 10, in which one master process (bottom) and four worker processes (top) are executed. The population is represented with larger squares, while the offspring is represented with smaller squares. The rectangular blocks in the region of the master process are executed sequentially, while the procedures performed in parallel are represented using two horizontal lines, and extend to the region corresponding to the workers. There are two stages in which the intensification procedure is applied to a set of individuals: after creating the initial population, and after creating the offspring. Those two blocks are the ones that are executed in parallel. By using the Message Passing Interface (MPI) [45], the individuals are distributed so that the worker processes carry out the intensification process and output the results. This distribution is carried out dynamically, meaning we wait for a process to finish carrying out the current intensification before sending it a new individual, continuing with this strategy as long as there are individuals remaining to be intensified. In the parallelized blocks, i.e., in the “Intensify”, the individuals that have already been sent to the work processes and were intensified are shown in green, those in which the intensification process is taking place are gray, and those that have not yet been intensified are red. The process will continue until the intensification has been performed on all the individuals, at which point the master process will continue with the next sequential blocks. The advantage of doing it dynamically is that it avoids imbalance problems when the tasks are heterogeneous. Such is the case with the local search, since the number of iterations to converge depends on the initial solution. Given that in the case of the LOP, the degree of heterogeneity is not high, in this research we selected population sizes that are multiples of the number of workers, which is an important consideration to enhance the efficiency. Finally, it is important to note that this way of parallelizing is independent of the components selected to carry out the intensification process, which is why the flowchart is presented in a general way. The parallelization shown was used to parallelize an extension of the MA-EDM algorithm presented in the next section.

### v2

@@ Abstract
The way heuristic optimizers are designed has evolved over the decades, as computing power has increased. Such has been the case for the Linear Ordering Problem (LOP), a field in which trajectory-based strategies led the way during the 1990s, but which have now been surpassed by memetic schemes. This paper focuses on understanding how the design of LOP optimizers will change in the future, as computing power continues to increase, yielding two main contributions. On the one hand, a metaheuristic was designed that is capable of effectively exploiting a large amount of computational resources, specifically, computing power equivalent to what a recent core can output during runs lasting over four months. Our analyses show that as the power of the computational resources increases, it will be necessary to boost the capacities of the intensification methods applied in the memetic algorithms to keep the population from stagnating. And on the other, the best-known results for today’s most challenging set of instances (xLOLIB2) were significantly outperformed. New bounds were established in this benchmark, which provides a new frame of reference for future research.
@@
An effective decision-making process usually involves solving highly complex optimization problems. The most general classification of optimization techniques distinguishes between exact and approximate optimization [1]. Given the complexity of many real-life optimization problems, exact optimization methods are usually not applicable within reasonable timeframes. Therefore, approximate methods have gained considerable popularity [2]. In this area, metaheuristics offer the possibility of increasing the generality of the developments, thus reducing the design efforts.
@@
One highly complex problem that has benefited greatly from the use of metaheuristics is the Linear Ordering Problem (LOP), which is an NP-hard combinatorial optimization problem with a large number of applications. The simplest definition of it is based on the matrix triangulation problem. Given a matrix , the problem is to determine a simultaneous permutation of the rows and columns of such that the sum of the entries above the main diagonal is maximized [3]. As a result, the problem can be formulated as the search for a permutation that maximizes (1):
@@
LOP has been applied in numerous areas, such as: determination of ancestry relationships in archaeology [4], straightline crossing minimization [5], and the development of probabilistic decision-making models [6]. Moreover, since many of these applications involve the use of large matrices, highly complex test sets have been defined. In particular, the most challenging validation set today is the LOLIB library [7], the latest version of which includes instances up to size 1000.
@@
Many diverse algorithms have been proposed to deal with the LOP [3]. Since there are no specific guides of steps to follow in the design of optimizers, researchers draw on the practical experience gained from similar problems [8]. In addition, the relative efficiency between metaheuristics depends largely on the computational capabilities available and the complexity of the instance to be solved, which accounts for the need for a wide variety of proposals [9]. The LOP is a fairly clear example of how computational resources and the complexity of instances affect the design of optimizers. Based on the historical review presented in [3], it is clear that effective optimizers have been evolving from ad-hoc heuristic methods, to trajectory metaheuristics, and on to the current population-based metaheuristics [10, 11].
@@
This paper falls within the field of approximate optimization and aims to contribute to the advancement of the design and development of efficient and robust optimizers. It focuses on the case of the LOP and seeks to answer the following research questions:
@@
Are the methods that generate the best solutions for the LOP able to do so in reasonable times, for example, under a day in the most challenging instances?
@@
Will today’s cutting-edge methods in the LOP continue to retain their prominence in the computing systems of the future, that is, after a significant increase in computing capabilities? Or, on the contrary, will they be unable to scale properly?
@@
Have optimal solutions already been achieved for the current most challenging datasets, or, on the contrary, is there still room for improvement?
In order to answer these questions at this time, it is necessary to make use of long-term executions and parallelizations to simulate the increased availability of computational resources. With this goal in mind, we conducted a long-term performance analysis of the current leading method for the LOP, the MA-EDM algorithm [11]. This analysis allowed us to identify some shortcomings in this method, and led us to conclude that, once the computing power is increased, MA-EDM has to be redesigned in order to leverage the additional computing power. In light of the weaknesses identified in the MA-EDM, a new optimizer called Memetic Algorithm with Explicit Diversity Management and Enhanced Intensification (MA-EDM) is proposed that applies an Iterated Local Search (ILS) as an improvement mechanism. As demonstrated in the experimental study, in high-performance environments, MA-EDM vastly outperforms the current best strategy. Moreover, by parallelizing MA-EDM, we were able to simulate scenarios with computing capabilities equivalent to those of a state-of-the-art core in executions lasting over 4 months. This experiment revealed that the best solutions found so far exhibit a significant gap with respect to the optimal ones, indicating that this set of instances has not yet been optimally solved, meaning, therefore, that more methodological advances are required. As an additional contribution, it should be noted that the new Best-Known Solutions (BKS) obtained for instances ranging in size from 300 to 1,000 will provide a frame of reference that will allow other researchers to better evaluate future advances.
The rest of the document is organized as follows. Section 2 discusses the design of approximate algorithms, with a special emphasis on Memetic Algorithms and their adaptations to the LOP. Section 3 focuses on describing the use of high-performance computing environments to achieve the objectives of this research. Section 4 justifies and presents MA-EDM. The experimental validation is presented in Section 5. Finally, the last section, discusses the conclusions and future areas of work.
@@ 2 Future and Limitations of Memetic Algorithms for the LOP
Evolutionary Algorithms (EAs) are a class of metaheuristics that draw their inspiration from biological evolution [12]. EAs are very effective at solving difficult optimization problems in a wide variety of fields [13]. However, they typically become more effective when combined with refinement mechanisms [14], combination known as Memetic Algorithms (MAs) [15]. In first-generation MAs, enhancement mechanisms were used statically to refine the solutions. Subsequently, a variety of more complex integration strategies were developed, ranging from schemes that consider a set of improvement mechanisms that are applied adaptively, to schemes that rely on the feedback provided by the improvement mechanisms to adapt internal parts of the EA. Despite the advances offered by recent MAs [16], their design, use and implementation is very complex due to the number of parameters and components they consider. In many problems, the first-generation MAs are still effective [17, 10], so both the simplest and the most complex methodologies continue to be investigated. Considering the number of problems in which MAs are the leading strategy [12], and that this is the case with the LOP, this paper focuses on these types of strategies. In particular, the MA that has yielded the current BKS for the LOP is a first-generation algorithm [11], so this is the main method used to conduct our analyses.
Since, for many problems, finding a local optimum is a computationally very expensive process, in such cases it is not feasible to use more complex schemes, which is why local search remains one of the most used methods in MAs; however, advances in hardware have also made it quite common for more complex trajectory metaheuristics and even mathematical programming strategies to be considered. Thus, there is a tendency to consider more complex refinement strategies, as the computation capabilities increase. In relation to the design of MAs, the problem of how often and how the improvement mechanism is used is a fundamental task that has been addressed in the literature in various ways [18]. Another crucial point is the definition of the neighborhood for the local search. In order to aid in the selection of the neighborhood, some heuristic procedures for analyzing the fitness landscape have been devised [15]. It is also important to note that in addition to the design issues implicit to MAs, as with any population-based algorithm, there are several additional characteristics that play an important role in the success of the optimization process. In particular, proper diversity management plays a critical role [19].
@@
In the case of MAs, problems with diversity management appear frequently. For example, in the RHGA algorithm [16], one of the most effective strategies for the TSP, great care had to be taken in selecting the frequency of application of the LKH improvement strategy to avoid premature convergence. This type of problem arises especially in the case of high dimensionality problems [20], so as the instances grow, it is quite common to have to change some design decisions. Issues related to instances of high dimensionality are not unique to MAs [21]. This problem is known as the “curse of dimensionality”, which occurs when an increase in search space size causes a significant loss of performance in optimization algorithms. The increase in the dimensionality of the instance causes the search space to grow exponentially in most problems, and as a result, the percentage of solutions evaluated decreases. However, this is not its only drawback, since it can, for example, cause premature convergence in subsets of variables [20]. Therefore, increasing the amount of computing power available (even exponentially) may not solve the challenges that arise when dealing with large instances. To address this problem, various techniques, such as dimensionality reduction techniques and coevolution, have been proposed [22].
This paper takes advantage of the demands associated with high dimensionality problems to further analyze and extend the metaheuristic schemes known to perform well with the LOP. To do this, we will work with instances of up to , that is, with a search space with approximately candidate solutions, and we will give the algorithm what it supposedly needs: more computing power. It is not at all foolhardy to think that in the near future, computing capacity will advance significantly, which will make more computing power available to algorithms. However, what we do not know is whether today’s methods will be able to continue to perform well when these resources are made available to them. This research provides an answer to this question.
@@ 3 Using High Performance Computing Environments
High performance computing (HPC) laboratories are typically used to reduce the time required to solve highly complex problems. Since the programming and bottlenecks of parallel algorithms differ from those of sequential algorithms, it is quite common for parallel algorithms to be developed that are significantly different from the direct parallelizations of the best sequential algorithms. Thus, in the case of Memetic Algorithms and population-based optimizers in general, various techniques have emerged to take advantage of this type of resource [23], such as master-worker parallelizations, island models, coevolution and cellular models. Among them, the only strategy that maintains the panmictic population-based model of the sequential algorithm is the master-worker model.
@@
The objective of using HPC in this research was not to take advantage of parallel computing resources to obtain high-quality results quickly, but to understand how the effectiveness of one of the leading algorithms in the area of LOP is affected by the increase in computing capabilities, as well as to propose modifications that allow the increase in computing power to be better utilized. Thus, the purpose of this work is not to innovate in parallelization models or find the best parallel models for the LOP, but to generate new knowledge in relation to the optimization of the LOP. In order to properly attain this objective, models should be used that keep the evolutionary process intact. Because of this, the synchronous master-worker model was selected. This scheme is standard in the parallelization of population-based algorithms and it based on the redistribution of a set of independent tasks among a set of processes that are exclusively dedicated to executing those tasks.
@@
The MA-EDM sequential algorithm is a standard first-generation MA. Its general behavior is described in Algorithm 1. A detailed description of each of its components is available at [11]. Each iteration operates with a population of individuals and consists of a binary tournament-based parent selection phase, a child-reproduction phase that considers the Cycle Crossover (CX), an improvement phase that applies a local search, and a replacement phase that uses the Best-Non Penalized strategy (BNP). Of all the above steps, the local search is the most computationally demanding process, since the neighborhood is considered based on shifts [24], so for instances of size , the step of ensuring that a local optimum has been found using incremental evaluation has a computational cost of while the rest of the components depend linearly or sublinearly on . Therefore, with large instances like the ones considered in this research, it is natural for the improvement process to be the task that is distributed in the parallelization. This distribution is done synchronously so as not to change the evolutionary process of the sequential scheme; that is, once a local search distribution point (or another improvement procedure) is reached, the master process halts until all the local searches are complete.
@@
The flowchart shown in Figure 1 illustrates the operation of the parallel version of MA-EDM (PMA-EDM). The diagram shows an example with a population of size 10, in which one master process (bottom) and four worker processes (top) are executed. The population is represented with larger squares, while the offspring is represented with smaller squares. The rectangular blocks in the region of the master process are executed sequentially, while the procedures performed in parallel are represented using two horizontal lines, and extend to the region corresponding to the workers. There are two stages in which the intensification procedure is applied to a set of individuals: after creating the initial population, and after creating the offspring. Those two blocks are the ones that are executed in parallel. By using the Message Passing Interface (MPI), the individuals are distributed so that the worker processes carry out the intensification process and output the results. This distribution is carried out dynamically, meaning we wait for a process to finish carrying out the current intensification before sending it a new individual, continuing with this strategy as long as there are individuals remaining to be intensified. In the parallelized blocks, i.e., in the “Intensify”, the individuals that have already been sent to the work processes and were intensified are shown in green, those in which the intensification process is taking place are gray, and those that have not yet been intensified are red. The process will continue until the intensification has been performed on all the individuals, at which point the master process will continue with the next sequential blocks. The advantage of doing it dynamically is that it avoids imbalance problems when the tasks are heterogeneous. Such is the case with the local search. Given that in the case of the LOP, the degree of heterogeneity is not high, in this research we selected population sizes that are multiples of the number of workers, which is an important consideration to enhance the efficiency. Finally, it is important to note that this way of parallelizing is independent of the components selected to carry out the intensification process, which is why the flowchart is presented in a general way. The parallelization shown was used to parallelize an extension of the MA-EDM algorithm presented in the next section.